OpenStudy (anonymous):

( Solve the equation 8cos0 +4=10. Express the answer in degree measure and approzimate the angular measure to the nearest tenth of a degree

OpenStudy (anonymous):

\[8\cos \theta + 4=10\]\[8\cos \theta = 6\]\[\cos \theta=.75\] Cosine is positive in the first and fourth quadrant. \[\theta =\cos^{-1} .75\]\[\theta =41.4 \] So 41.4 degrees or 318.6 degrees